You have just made your first $4,500 contribution to your individual retirement account. Assume you earn a 10.55 percent rate of return and make no additional contributions.

Requirement 1:

What will your account be worth when you retire in 44 years? (Enter rounded answer as directed, but do not use rounded numbers in intermediate calculations. Round your answer to 2 decimal places (e.g., 32.16).)

  Amount $   
Requirement 2:

What if you wait 10 years before contributing? (Enter rounded answer as directed, but do not use rounded numbers in intermediate calculations. Round your answer to 2 decimal places (e.g., 32.16).)

  Amount $   
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1.A=4500*(1.1055)^44

=$371,361.93(Approx).

2.A=4500*(1.1055)^34

=$136,210.29(Approx).
Answer #2

 Calculation :-

Requirement 1: Growing for 44 Years

You invest $4,500 today at an annual return of 10.55%. The formula for future value is:

Future Value=Present Value×(1+Interest Rate)Time

Plugging in the numbers:

FV=4500×(1+0.1055)44FV=4500×(1.1055)44

Using a calculator:

  1. 1.10554489.49521

  2. Multiply by $4,500:

4500×89.49521=402,728.44

Amount ≈ $402,728.44



Requirement 2: Waiting 10 Years, Then Investing for 34 Years

If you wait 10 years, you only have 34 years left to grow the money.

Using the same formula:

FV=4500×(1.1055)34

Calculating:

  1. 1.10553432.54369

  2. Multiply by $4,500:

4500×32.54369=146,446.60

Amount ≈ $146,446.60


answered by: Harshwardhan kunal